Menthol (M.M=156.3g/mol),a strong smelling substance used in cough drops is a compound of carbon , hydrogen , and oxygen. When 0.1595g of menthol was subjected to combustion analysis, it produced 0.449g of CO2 and 0.184g of H2O...what is a menthol's molecular formula ?
1
Expert's answer
2017-03-29T07:29:18-0400
Answer
CxHyOz
CO₂->C
0.449g/44g/mol=xg(C)/12g/mol;m(C)=0.1225g
H₂O->2H
o.184g/18g/mol=yg(H)/2g/mol;m(H)=0.0204g
m(O)=0.1595-(0.1225+0.0204)=0.0166g
x:y:z=0.1225/12:0.0204/1:0.0166/16
x:y:z=10.2:20.4:1
x=10;y=20;z=1
M(C₁₀H₂₀O)=156.3g/mol
C₁₀H₂₀O